OpenStudy (anonymous):

what are the solutions of the system y=x^2-6x+7 and y=-x+13

hartnn (hartnn):

equate them x^2-6x+7 = -x+13 this is a quadratic in x can you find 2 values of x from this ?

OpenStudy (anonymous):

a.) (-1,14) and (6,7) b.) (-1,7) and (6,14) c.) (-1,13) and (6,7) d.) no solutions I am leaning more towards C or D?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

did u try to solve that quadratic ?

OpenStudy (anonymous):

Uhh I don't really know how to solve quadratic

hartnn (hartnn):

x^2-6x+7 = -x+13 x^2 -6x +x +7 -13 = 0 x^2 -5x -6 = 0 can u tell me 2 numbers whose sum is -5 and product = -6

OpenStudy (anonymous):

3 and 2

OpenStudy (anonymous):

well hmm wait that doesn't work because uh a - times a - in a + but -3 +2 will not be -5

hartnn (hartnn):

yeah try once more :)

OpenStudy (anonymous):

-6 and 1

hartnn (hartnn):

YES!

hartnn (hartnn):

so, i split -5x as -6x+x

hartnn (hartnn):

x^2 -6x +x -6 = 0 x (x-6) +1 (x-6) = 0 (x-6) (x+1) = 0 got this ?? can u proceed ?

OpenStudy (anonymous):

I don't think you can continue I mean yeah x=-6 and x=1 but that answer isn't showing here so will it be no solution?

hartnn (hartnn):

not actually x-6 = 0 gives you x= 6 x+1 = 0 gives you x = -1

hartnn (hartnn):

so, find the value of y when x=6 y = -x +13 y = -6 + 13 =.. ?

OpenStudy (anonymous):

Ahh so it is opposite...I think I get it now. and y= -6+13=0?

hartnn (hartnn):

-6+13 is +7 so one of your point is 6,7

OpenStudy (anonymous):

ahh okay...and how will i solve for the other point with the -1?

hartnn (hartnn):

y = -x +13 y = -(-1) + 13 y = 1+13 = 14 so other point = -1,14

OpenStudy (anonymous):

Woo thanks so much :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Can I have more help?

hartnn (hartnn):

sure, just ask new question in new post :)